Walton Company manufactures two products. The budgeted per-unit contribution margin for each product follows:

Super Supreme s 97 $127 Sales price Variable cost per unit Contribution margin per unit (66(82) $31 45

Walton expects to incur annual fixed costs of $133,760. The relative sales mix of the products is 70 percent for Super and 30 percent for Supreme.

Required

  1. Determine the total number of products (units of Super and Supreme combined) Walton must sell to break even.

  2. How many units each of Super and Supreme must Walton sell to break even?

Required a. Determine the total number of products (units of Super and Supreme combined) Walton must sell to break even. b. How many units each of Super and Supreme must Walton sell to break even? (For all requirements, do not round intermediate calculations.) a. Total number of products b. Product Super units units units Product Supreme
Answer #1
Particulars Super Supreme
A. Contribution margin 31 45
B. Sales mix 70% 30%
D.Weighted Contribution Margin Per Unit (Contribution margin per unit *Sales Mix) 21.7 13.5

Part-a)

Total Units to Break Even = Fixed Cost / Total Weighted Contribution Margin Per Unit

=$133,760/(21.7+13.5)

=3,800 units

Part-b)

No of units of Super sold =3800*70%=2,660

No of units of Supreme sold =3800*30%=1,140

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY GIVING A ??

THANK YOU